# Closed-Form Approximation of the Total Variation Proximal Operator

### v1

@@ 6 Appendix
$=\bar{h}({\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+\frac{1}{\tau}({\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}}))^{{\mathsf{T}}}({\bm{y}}-{\mathcal{S}}_{\tau}({\bm{z}}))\;,$ (21)
where the last inequality comes from the definition of adjoint $({\bm{W}}{\bm{x}}_{1})^{{\mathsf{T}}}{\bm{x}}_{2}={\bm{x}}_{1}^{{\mathsf{T}}}({\bm{W}}^{{\mathsf{T}}}{\bm{x}}_{2})$, the fact that ${\bm{W}}^{{\mathsf{T}}}{\bm{W}}={\bm{I}}$, and that $\|{\bm{W}}^{T}\|_{2}=1$. In the last line, we used the fact that ${\bm{u}}={\mathsf{prox}}_{\tau\bar{h}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ from Eq. 15 and ${\mathcal{S}}_{\tau}({\bm{z}})={\bm{W}}^{{\mathsf{T}}}{\bm{u}}$
Figure 5: A visual illustration of the characterization of the projection as ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}$. For any ${\bm{z}}\in\mathbb{R}^{n}$, the vector ${\bm{W}}{\bm{z}}\in{\mathcal{U}}:=\{{\bm{u}}|\ {\bm{u}}={\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\}$. Applying the soft-thresholding function gives ${\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})$, not necessarily in ${\mathcal{U}}$. Applying ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ projects back to ${\mathcal{U}}$, which can be represented as ${\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}$.
@@
${\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}\;.$ (22)
That is, ${\bm{\delta}}$ is the orthogonal component of ${\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ to the subspace ${\mathcal{U}}:=\{{\bm{u}}|\ {\bm{u}}={\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\}$. Figure 5 provides a visual representation.
For both the anisotropic and isotropic versions, the soft-thresholding operator ${\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ at most changes each of the $nd$ difference elements of ${\bm{W}}{\bm{z}}$ by $2\tau\sqrt{d}$, giving the bound
$\|{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\|_{2}$ $\leq\sqrt{nd(2\tau\sqrt{d})^{2}}$
$=\sqrt{4\tau^{2}nd^{2}}$
$=2\tau d\sqrt{n}\;.$ (23)
Since the projection ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ minimizes the $\ell_{2}$ distance between ${\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ and ${\mathcal{U}}$, the following inequality holds
$\|{\bm{\delta}}\|_{2}$ $=\mathop{\mathsf{min}}_{{\bm{w}}\in{\mathcal{U}}}\|{\bm{w}}-{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})\|_{2}$
$\leq\|{\bm{w}}-{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})\|_{2}\qquad\forall{\bm{w}}\in{\mathcal{U}}\;.$
@@
$\|{\bm{\delta}}\|_{2}\leq\|{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\|_{2}\leq 2\tau d\sqrt{n}\;.$
Furthermore,
$\|{\bm{\delta}}^{\text{dif}}\|_{2,1}\leq\|{\bm{\delta}}^{\text{dif}}\|_{1,1}$ $=\|{\bm{\delta}}^{\text{dif}}\|_{1}$
$\leq\sqrt{nd}\|{\bm{\delta}}^{\text{dif}}\|_{2}$
$\leq\sqrt{nd}\|{\bm{\delta}}\|_{2}$
@@
$=2\sqrt{d}\|[{\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}$
$=2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}]^{\text{dif}}\|_{p,1}$
$\leq 2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}+2\sqrt{d}\|{\bm{\delta}}^{\text{dif}}\|_{p,1}$
$\leq 2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}+(2\sqrt{d})(2\tau nd^{3/2})$
$=\bar{h}({\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+4\tau nd^{2}\;.$
@@
$\bar{h}({\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}))\geq{h}({\mathcal{S}}_{\tau}({\bm{z}}))-4\tau nd^{2}\;.$ (24)
@@
${h}({\bm{y}})\geq\bar{h}({\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+\frac{1}{\tau}({\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}}))^{{\mathsf{T}}}({\bm{y}}-{\mathcal{S}}_{\tau}({\bm{z}}))\;.$
@@
Corollary 1.
For any $\tau>0$ and for any ${\bm{z}}\in\mathbb{R}^{n}$
$\big{\|}{\mathsf{prox}}_{\tau{h}}({\bm{z}})-{\mathcal{S}}_{\tau}({\bm{z}})\big{\|}_{2}\leq\tau\epsilon_{1}\;.$
Proof.
Define function $\Phi_{\tau}$ for ${\bm{z}}\in\mathbb{R}^{n}$ as
$\Phi_{\tau}({\bm{x}})=\frac{1}{2\tau}\|{\bm{x}}-{\bm{z}}\|_{2}^{2}+\tau{h}({\bm{x}})\;.$
By using Proposition 2.5 (2.) of [48] with this paper’s Proposition 2 (b), we have
${\mathcal{S}}_{\tau}({\bm{z}})={\mathsf{prox}}_{\tau{h}}({\bm{z}}+{\bm{\delta}})\quad\text{ where}\quad\|{\bm{\delta}}\|_{2}\leq\tau\epsilon_{1}$
$\implies$
$0\in\partial_{(\tau\epsilon_{1})^{2}/(2\tau)}\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))\;.$
Using [43, 44] implies that
$\big{\|}{\mathsf{prox}}_{\tau{h}}({\bm{z}})-{\mathcal{S}}_{\tau}({\bm{z}})\big{\|}_{2}\leq\tau\epsilon_{1}\;.$
@@
This lemma makes use of the following subdifferential sum property: for any $f_{1},\dots,f_{m}$ real valued convex functions
@@
First consider the anisotropic version
@@
Consider $\partial\big{(}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{1}\big{)}$, whose norm is bounded by $\sqrt{d}$. By Eq. (25), letting ${\bm{g}}\in\partial\bar{h}^{\text{a}}({\bm{u}})$, we can bound
@@
Consider $\partial\big{(}\|[{\bm{u}}^{\text{dif}}]_{n}\|_{2}\big{)}$, whose norm is bounded by $1$. By Eq. (25), letting ${\bm{g}}\in\partial\bar{h}^{\text{i}}({\bm{u}})$, we obtain the bound

### v2

@@ 6 Appendix
$=\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+\frac{1}{\tau}({\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}}))^{{\mathsf{T}}}({\bm{y}}-{\mathcal{S}}_{\tau}({\bm{z}}))\;,$ (26)
where the last inequality comes from the definition of adjoint $({\bm{W}}{\bm{x}}_{1})^{{\mathsf{T}}}{\bm{x}}_{2}={\bm{x}}_{1}^{{\mathsf{T}}}({\bm{W}}^{{\mathsf{T}}}{\bm{x}}_{2})$, the fact that ${\bm{W}}^{{\mathsf{T}}}{\bm{W}}={\bm{I}}$, and that $\|{\bm{W}}^{T}\|_{2}=1$. In the last line, we used the fact that ${\bm{u}}={\mathsf{prox}}_{\tau\bar{h}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ from Eq. 19 and ${\mathcal{S}}_{\tau}({\bm{z}})={\bm{W}}^{{\mathsf{T}}}{\bm{u}}$ .
Figure 5: A visual illustration of the characterization of the projection as ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}$. For any ${\bm{z}}\in\mathbb{R}^{n}$, the vector ${\bm{W}}{\bm{z}}\in{\mathcal{U}}:=\{{\bm{u}}|\ {\bm{u}}={\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\}$. Applying the soft-thresholding function gives ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$, not necessarily in $\mathcal{U}$}. Applying ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ projects back to $\mathcal{U}$}, which can be represented as ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}$.
@@
${\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\beta}}\;.$ (27)
That is, $\bm{\beta}$} is the orthogonal component of ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ to the subspace ${\mathcal{U}}:=\{{\bm{u}}|\ {\bm{u}}={\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\}$. Figure 5 provides a visual representation.
Since the projection ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ minimizes the $\ell_{2}$ distance between ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ and $\mathcal{U}$}, the following inequality holds
$\|{\bm{\beta}}\|_{2}$ $=\mathop{\mathsf{min}}_{{\bm{w}}\in{\mathcal{U}}}\|{\bm{w}}-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})\|_{2}$
$\leq\|{\bm{w}}-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})\|_{2}\qquad\forall{\bm{w}}\in{\mathcal{U}}\;.$
@@
$\|{\bm{\beta}}\|_{2}\leq\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\|_{2}\leq 2\tau d\sqrt{n}\;,$
where the last inequality comes from (23). Letting ${\bm{\beta}}^{\text{dif}}$ be the components of $\bm{\delta}$} corresponding to difference coefficients,
$\|{\bm{\beta}}^{\text{dif}}\|_{2,1}\leq\|{\bm{\beta}}^{\text{dif}}\|_{1,1}$ $=\|{\bm{\beta}}^{\text{dif}}\|_{1}$
$\leq\sqrt{nd}\|{\bm{\beta}}^{\text{dif}}\|_{2}$
$\leq\sqrt{nd}\|{\bm{\beta}}\|_{2}$
@@
$=2\sqrt{d}\|[{\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}$
$=2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\beta}}]^{\text{dif}}\|_{p,1}$
$\leq 2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}$
$\qquad\qquad\qquad+2\sqrt{d}\|{\bm{\beta}}^{\text{dif}}\|_{p,1}$
$\leq 2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}$
$\qquad\qquad\qquad+(2\sqrt{d})(2\tau nd^{3/2})$
$=\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+4\tau nd^{2}\;.$
@@
$\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))\geq{h}({\mathcal{S}}_{\tau}({\bm{z}}))-4\tau nd^{2}\;.$ (28)
@@
${h}({\bm{y}})\geq\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+\frac{1}{\tau}({\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}}))^{{\mathsf{T}}}({\bm{y}}-{\mathcal{S}}_{\tau}({\bm{z}}))\;.$
@@
This lemma makes use of the following subdifferential sum property: for any functions $f_{1},\dots,f_{m}\in\Gamma^{0}$ such that the sets $\text{ri}(\text{dom}f_{i})$ have a point in common, where ri is the relative interior, then
@@
See Theorem 23.8 in [66] and Corollary 16.48 in [70]. First consider the anisotropic version
@@
Since the domain of each $||\cdot||_{1}$ is all of $\mathbb{R}^{2nd}$, the relative interiors of the domains have a nonempty intersection, and so we can apply Eq. (29). Consider $\partial\big{(}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{1}\big{)}$, whose norm is bounded by $\sqrt{d}$. By Eq. (29), letting ${\bm{g}}\in\partial\bar{h}^{\text{a}}({\bm{u}})$, we can bound
@@
Similarly, the domain of each $||\cdot||_{2}$ is all of $\mathbb{R}^{2nd}$, and so we can apply Eq. (29). Consider $\partial\big{(}\|[{\bm{u}}^{\text{dif}}]_{n}\|_{2}\big{)}$, whose norm is bounded by $1$. By Eq. (29), letting ${\bm{g}}\in\partial\bar{h}^{\text{i}}({\bm{u}})$, we obtain